OpenStudy (ria23):

state the Domain and range~ k(x)=4x+3/x^2-1

OpenStudy (ria23):

Domain= \[\left\{ x:x \mathbb{R}-1\le x \le1 \right\}\] I'm not sure about range... @mathmath333

OpenStudy (mathmath333):

the degree is different on numerator and denominator so it will be \(\large\tt \begin{align} \color{black }{k(x)\in \mathbb{R} }\end{align}\)

OpenStudy (ria23):

This will be like the other one... The uhm.... \[a^{3}-b^{3}=(a-b)(a^{2}+ab+b^{2})\]?

OpenStudy (mathmath333):

yes so the domain will be \(\Large x\neq 1\\ \Large x\neq -1\)

OpenStudy (mathmath333):

and its \(\Large x^2-1=(x+1)(x-1)\)

OpenStudy (mathmath333):

its not the power \(\Large 3\)

OpenStudy (ria23):

So I would write that, but with the K(x)ER in front?

OpenStudy (mathmath333):

it depends how ur teacher taught u

OpenStudy (ria23):

I don't think it matters. So it can be written\[k(x)\in \mathbb{R} x^{2}-1(x+1)(x-1)\]? Or without the x^2-1 in the front?

OpenStudy (mathmath333):

idk i would write it as domain= \(\large\tt \begin{align} \color{black}{\{x:x\in \mathbb{R} :x\neq 1 \land x\neq-1}\}\end{align}\)

OpenStudy (ria23):

Can yhu show me the steps to find the range? I couldn't get that one. ;c

OpenStudy (mathmath333):

find the domain of inverse of function their domain will be the range